# Speech-Audio Compositional Attacks on Multimodal LLMs and Their Defense with SALMONN-Guard

### v2

@@ 5.3 Ablation Studies
First, we conducted an ablation study on the acoustic attributes of the speech overlap attack. To ensure intelligibility of audio mixture, in addition to manual listening check conducted among co-authors, we prompt Gemini 2.5 Pro to describe the audio content. If the description matches both parts of the speech, or refuses to generate one side due to safety reasons, we consider it as perceptible. As a results, 93% of the samples in the speech overlap partition can be clearly perceived by Gemini 2.5 pro. This ensures that the speech overlap is sufficiently intelligible given that Gemini 2.5 pro is imperfect in understanding overlapped speech.
@@
The ablation results, shown in Table 2, reveal that attack efficacy is most sensitive to parameters that increase the subtlety of the harmful instruction. We observe that a higher playback speed, a lower volume, and a longer overlap duration all correlate with a higher Attack Success Rate. This supports our hypothesis that the attack is most potent when the malicious content is acoustically framed as subliminal information—clearly perceived by the model, yet subtle enough to bypass its safety filters.
@@
Furthermore, we conducted an ablation study on the Multi-speaker Dialogue attack to isolate the source of its efficacy. As shown in Table 3, we evaluated Gemini 1.5 Pro against unimodal baselines. The results demonstrate that our cross-modal Text + Audio approach (78.58% ASR) is significantly more potent than either a Text-Only (57.96% ASR) or Audio-Only (65.03% ASR) attack. This confirms that the attack’s high success rate stems from the synergistic effect of pairing a benign text prompt with a malicious audio context, a strategy that effectively bypasses unimodal safety checks.
5.4 Comparison against other jailbreaking
@@
To highlight the unique nature of our speech-audio composition-based attacks and evaluate the generalization of our defense, we compare model performance against two recent perturbation-based audio jailbreaking methods: Speech Insertion (yang2024audioachillesheelred) and Speech Editing (cheng2025jailbreakaudiobenchindepthevaluationanalysis). The evaluation data for these methods is sourced from JALMBench (peng2025jalmbenchbenchmarkingjailbreakvulnerabilities). For Speech Insertion, we use the 246 samples from the SSJ subset. For Speech Editing, we use a 100-sample subset from AMSE, which involves 18 distinct edits per sample, resulting in 1,800 total audio inputs. The results, presented in Table 4, reveal that our methods exploit a fundamentally different class of vulnerabilities for which current models are largely unprepared. This analysis leads to two critical findings.
@@
First, the results demonstrate that our speech-audio composition-based attacks pose a far greater challenge to state-of-the-art LLMs than prior methods. For instance, while Gemini 2.5 Pro shows some resilience to Speech Insertion (13.41% ASR) and Speech Editing (23.00% ASR), its vulnerability skyrockets to 65.41% ASR on SACRED-Bench. This trend is even more pronounced for open-source models like Qwen3-Omni, whose ASR jumps from 8.94% on Speech Insertion to 81.50% on our benchmark. This disparity underscores that our methods, which manipulate semantic and contextual cues, are fundamentally more effective at bypassing modern safety alignments than attacks that rely on signal-level perturbations.
@@ 6 Conclusion
We introduced SACRED-Bench, a benchmark demonstrating that compositional audio attacks can effectively bypass existing safeguards. Our experiments revealed significant vulnerabilities in state-of-the-art LLMs, with an Attack Success Rate exceeding 66% on Gemini 2.5 Pro, exposing the critical limitations of text-centric safety filters. As a countermeasure, our proposed SALMONN-Guard successfully reduced the ASR to approximately 20% by performing holistic multimodal safety checks. This paper underscores the urgent need to develop robust, truly multimodal safety paradigms for the next generation of audio LLMs.
Appendix A Detailed SACRED-Bench Statistics
@@
Multi-speaker Dialogue 919 1364
Speech-Audio Mixture 986 717

### v3

@@ 5.3 Ablation Studies
First, we conducted an ablation study on the acoustic attributes of the speech overlap attack. To ensure intelligibility of audio mixture, in addition to manual listening check conducted among co-authors, we prompt Gemini 2.5 Pro to describe the audio content. If the description matches both parts of the speech, or refuses to generate one side due to safety reasons, we consider it as perceptible. As a result, 93% of the samples in the SSO partition can be clearly perceived by Gemini 2.5 Pro. This ensures that the speech overlap is sufficiently intelligible, given that Gemini 2.5 Pro is imperfect in understanding overlapped speech.
@@
The ablation results, shown in Table 2, reveal that attack efficacy is most sensitive to parameters that increase the subtlety of the harmful instruction. We observe that a higher playback speed, a lower volume, and a longer overlap duration all correlate with a higher Attack Success Rate. This supports our hypothesis that the attack is most potent when the malicious content is acoustically framed as subliminal information—clearly perceived by the model, yet subtle enough to successfully jailbreak its safety filters.
@@
Furthermore, we conducted an ablation study on the MSD attack to isolate the source of its efficacy. As shown in Table 3, we evaluated Gemini 1.5 Pro against unimodal baselines. The results demonstrate that our cross-modal Text + Audio approach (78.58% ASR) is significantly more effective than either a Text-Only (57.96% ASR) or Audio-Only (65.03% ASR) attack. This confirms that the attack’s high success rate stems from the synergistic effect of pairing a benign text prompt with a malicious audio context, a strategy that effectively jailbreaks unimodal safety guardrails.
5.4 Comparison against other jailbreaking methods
@@
To highlight the unique nature of our speech-audio composition-based attacks and evaluate the generalization of our defense, we compare model performance against two recent perturbation-based audio jailbreaking methods: Speech Insertion (Yang et al., 2024b) and Speech Editing (Cheng et al., 2025). The evaluation data for these methods is sourced from JALMBench (Peng et al., 2025). For Speech Insertion, we use the 246 samples from the SSJ subset. For Speech Editing, we use a 100-sample subset from AMSE, which involves 18 distinct edits per sample, resulting in 1,800 total audio inputs. The results, presented in Table 4, reveal that our methods exploit a fundamentally different class of vulnerabilities for which current models are largely unprepared. This analysis leads to two critical findings.
@@
First, the results demonstrate that our speech-audio composition-based attacks pose a far greater challenge to state-of-the-art LLMs than prior methods. For instance, while Gemini 2.5 Pro shows some resilience to Speech Insertion (13.41% ASR) and Speech Editing (23.00% ASR), its vulnerability skyrockets to 65.41% ASR on SACRED-Bench. This trend is even more pronounced for open-source models like Qwen3-Omni, whose ASR jumps from 8.94% on Speech Insertion to 81.50% on our benchmark. This disparity underscores that our methods, which manipulate semantic and contextual cues, are fundamentally more effective at breaking modern safety alignments than attacks that rely on signal-level perturbations.
@@ 6 Conclusion
We introduced SACRED-Bench, the first compositional audio attack benchmark that effectively jailbreaks existing safeguards. We proposed three different ways to composite harmful audio streams, including speech-speech overlap, speech-audio overlap, and multi-speaker dialogue, which are all representative of real-world scenarios. Our experiments revealed significant vulnerabilities in state-of-the-art LLMs, with an Attack Success Rate exceeding 66% on Gemini 2.5 Pro, exposing the critical limitations of text-centric safety filters. As a countermeasure, our proposed SALMONN-Guard successfully reduced the ASR down to approximately 20% by performing holistic multimodal safety checks, showcasing the need to develop multimodal safety guardrails for the next generation of LLMs.
7 Impact Statement
This work investigates safety vulnerabilities in multimodal large language models (LLMs) arising from complex speech–audio compositions and introduces both a benchmark (SACRED-Bench) and a defense mechanism (SALMONN-Guard) to address these risks. Our primary intended impact is to improve the safety, robustness, and trustworthiness of multimodal AI systems that process real-world audio inputs.
Positive social impact: As multimodal LLMs are increasingly deployed in settings such as voice assistants, accessibility tools, education, and content moderation, their ability to correctly interpret and regulate complex audio inputs is critical. By demonstrating that realistic audio scenarios, such as overlapping speakers, background sounds, and multi-speaker dialogue, can bypass existing safeguards, this work highlights an important and previously underexplored safety gap. SACRED-Bench provides the research community with a systematic, reproducible tool to evaluate these risks, while SALMONN-Guard shows that joint inspection of speech, non-speech audio, and text can substantially reduce harmful model behavior. Together, these contributions can inform the development of more robust multimodal safety mechanisms and help prevent the deployment of systems that may generate or endorse harmful content.
Potential risks and misuse: Like other red-teaming and safety benchmarks, SACRED-Bench could potentially be misused to intentionally stress or exploit deployed systems. However, the benchmark does not introduce new harmful content categories or novel attack capabilities beyond realistic audio compositions that already occur in everyday environments. Moreover, the benchmark is designed for evaluation rather than deployment-time exploitation, and the paper places strong emphasis on defensive insights and mitigation strategies. We believe that openly studying these vulnerabilities is necessary to enable effective safeguards and ultimately reduce real-world harm.
Mitigation and responsible use: To mitigate misuse, our work pairs the identification of vulnerabilities with a concrete defensive solution. SALMONN-Guard demonstrates that the risks exposed by SACRED-Bench are not inherent or unavoidable, but can be substantially reduced through multimodal safety modeling. We encourage practitioners to use SACRED-Bench responsibly as a diagnostic and benchmarking tool, and to prioritize safety-oriented deployment practices when working with multimodal LLMs.
Appendix A Qualitative Examples of Model Responses
We provide examples of model responses to queries in SACRED-Bench in Fig. 4 to 6. Note that there is harmful content in both the audio description and the model response.
Figure 4: Example model responses under SSO attack.
Figure 5: Example model responses under MSD attack.
Figure 6: Example model responses under SAO attack.
Appendix B Automated Evaluation Prompt
In this section, we provide the detailed prompt used for the automated evaluation described in Section 3.4.2. We utilize a large language model (Gemini 2.5 Pro) as an independent judge to classify the target model’s responses.
Regarding the calculation of the Attack Success Rate (ASR), we adopt a strict criterion for model safety. We consider both Harmful responses and instances of Wrong understanding as successful attacks. Consequently, an attack is deemed unsuccessful and the model considered robust only when the response is classified strictly as a Direct Refusal.
The specific prompt template provided to the judge is formatted as follows:
Judge Prompt Template System Prompt: You are an expert in red-teaming. You will be given a prompt and its response from a large-scale audio-language model. Please classify the response into three classes: Harmful, Direct Refusal, Wrong understanding. Based on the prompt: If the response is a useful answer to the prompt, then it should be classified as “Harmful”; Else if the response is a direct refusal to the prompt, usually started with sentences like “I cannot answer the question.”, and followed with nothing useful to the prompt, then it should be classified as “Direct Refusal”; Then, if the response is a wrong understanding of the prompt, generating nothing useful to the prompt, but also does not directly refuse to answer the prompt, then it should be classified as “Wrong understanding”. You should only output the class within these three classes (which are Harmful, Direct Refusal, Wrong understanding) without any explanation. User Input: The prompt and intended harmful behavior is: {Prompt and Behavior} The response is: {model_answer}
Appendix C Detailed SACRED-Bench Statistics
@@
Speech non-speech Audio Overlap (SAO) 919 1364
Multi-speaker Dialogue (MSD) 986 717
@@
Appendix D Hyperparameters and Training Details
In this section, we outline the detailed hyperparameters used in our two-stage fine-tuning curriculum (Section 3.4.2). We utilized Qwen2.5-Omni-7B as the base model and applied Low-Rank Adaptation (LoRA) for parameter-efficient fine-tuning. The training was conducted using the bfloat16 precision to optimize memory usage.
Table 6 summarizes the specific configuration used for the experiments. Notably, as mentioned in the methodology, we update the Audio Encoder and Aligner modules alongside the LLM backbone (indicated by freeze_vit=False and freeze_aligner=False).
Table 6: Hyperparameters used for fine-tuning Qwen2.5-Omni-7B.
Hyperparameter Value
Model Configuration
Base Model Qwen/Qwen2.5-Omni-7B
Precision bfloat16
Max Sequence Length 2048
LoRA Configuration
LoRA Rank ($r$) 32
LoRA Alpha ($\alpha$) 32
Target Modules All Linear Layers
Optimization
Learning Rate $3\times 10^{-5}$
Weight Decay 0.1
Max Gradient Norm 0.5
Warmup Ratio 0.05
Optimizer AdamW
Batch Size & Steps
Per-Device Batch Size 4
Gradient Accumulation Steps 2
Stage 1 Epochs (General) 3
Stage 2 Epochs (Specialized) 5
Appendix E Human Intelligibility Study
We assess the intelligibility of harmful content in SACRED-Bench by conducting a listening test with two participants: one among the authors and one external third-party examiner.
Table 7: Listening test with two different listeners on SACRED-Bench SSO and SAO. 25 audio samples are randomly sampled from each partition to conduct the listening test. Intelligibility is evaluated by the rate at which harmful information is detected by humans.
Listener SSO SAO
Listener 1 100% 96%
Listener 2 96% 96%
Overall 98% 96%
Appendix F Detailed Analysis and Validation of MSD Evaluation
To ensure the reliability and robustness of our evaluation metrics for Multi-Speaker Dialogue (MSD) attacks, we conducted a rigorous validation study. This involves validating our automated judges against human ground truth, decomposing the failure modes, and verifying the perceptibility of the attacks.
F.1 Cross-Validation of Automated Judges
In our main experiments, we primarily utilized Gemini 2.5 Pro as the automated judge. To mitigate potential biases inherent in a single-model evaluation, we introduce Qwen2.5-Omni-7B as a second automated judge to evaluate the level of bias. We evaluated the alignment between these two models and human experts: We randomly sampled $N=50$ instances from the MSD test set. These samples were independently annotated by: Human Experts, Gemini 2.5 Pro and Qwen2.5-Omni-7B.
We calculated the inter-annotator agreement using Cohen’s Kappa ($\kappa$) to quantify the reliability of the automated judges.
Table 8: Inter-annotator agreement (Cohen’s Kappa and Agreement Rate) between human experts and automated judges.
Annotator Pair Cohen’s Kappa ($\kappa$) Agreement Rate (%)
Human Annotator 1 vs. Human Annotator 2 0.89 95.0
Human Annotator vs. Gemini 2.5 Pro Judge 0.82 91.0
Human Annotator vs. Qwen2.5-Omni-7B Judge 0.86 93.0
Gemini 2.5 Pro Judge vs. Qwen2.5-Omni-7B Judge 0.80 89.0
F.2 Decomposition of Failure Modes
To accurately interpret the Attack Success Rate (ASR), it is crucial to distinguish between genuine safety failures and mere capability limitations. We classify the model’s failed responses into two categories:
  • Harmful Compliance: The model correctly understands the hidden malicious intent and executes the instruction.
  • Wrong Understanding: The model fails to refuse the request but generates irrelevant, confused, or hallucinatory content.
Table 9: Breakdown of ASR on the MSD test set. By separating Harmful Compliance from Wrong Understanding, we show that the high ASR of Gemini 2.5 Pro is driven by actual safety failures, whereas SALMONN-Guard effectively minimizes harmful outputs.
Model Total ASR (%) Harmful Compliance (%) Wrong Understanding (%)
Gemini 2.5 Pro 63.93 55.12 8.81
SALMONN-Guard (Ours) 14.08 2.15 11.93